OpenStudy (teller):

Can someone tell me if this is correct? The half-life of a radioactive isotope is the time it takes for a quantity of the isotope to be reduced to half its initial mass. Starting with 215 grams of a radioactive isotope, how much will be left after 4 half-lives?Use the calculator provided and round your answer to the nearest gram. f=215(1+0.005)^4=219.33235763 so would it be 219.32?

OpenStudy (teller):

@phi can you see if I got this right?

OpenStudy (phi):

The half-life of a radioactive isotope is the time it takes for a quantity of the isotope to be reduced to half its initial mass that means if you start with 100, after one half-life, you will have 50 your answer should be smaller than the amount you start with.

OpenStudy (teller):

Got a little confused there but would it be 210

OpenStudy (phi):

where did you get your formula? I would use 215*(0.5)^n where n is the number of half-lifes

OpenStudy (teller):

i got it from the explain tab it says to use those formulas so would it be 13.4375

OpenStudy (phi):

yes, that looks better you could do it like this 215 --> 107.5 (1 half life) 107.5/2 = 53.75 (2nd) 53.75/2 = 26.875 (3rd) 26.875/2= 13.4375 (4 half lifes) the formula is faster to use.

OpenStudy (teller):

but it said to the nearest gram so would it be 13.42?

OpenStudy (phi):

the starting amount is in grams so the answer 13.4375 is in grams round to the nearest gram means the nearest whole number you have 13.4 (plus a tiny bit more) is 13.4 closer to 13 or to 14 ?

OpenStudy (teller):

13

OpenStudy (teller):

so the answer would be 13?

OpenStudy (phi):

yes 13 the rule in rounding is look at the next lower digit 13.4 <-- look at the 4 if 5 or bigger round up (to 14) else round down to 13

OpenStudy (teller):

Thank you mind helping me with another one ?Once I get it solved since this is my last question.